Circular No. 8555

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005cs IN M51 M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center

for Astrophysics, report that a spectrogram (range 340–740 nm) of SN 2005cs (cf. IAUC 8553), obtained by R. Hutchins on June 30.23 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a young type-II supernova. The spectrum consists of a blue continuum and P-Cyg profiles of the Balmer and He lines. Adopting the NED recession velocity of 600 km/s for the host galaxy (from Brian et al. 1996, Ap.J. 473, 130), the expansion velocity derived from the minimum of the Hβ line is ∼ 7500 km/s. Zero-velocity interstellar Na I D absorption with an equivalent width of ∼ 0.02 nm is detected in the spectrum of SN 2005cs, indicating gas along the line-of-sight in our galaxy and thus suggesting some reddening by dust. Schlegel et al. (1998, Ap.J. 500, 525) estimate E(B − V) = 0.035 mag of Galactic reddening along the line-of-sight to M51. Interstellar Na I D absorption is detected also at the redshift of the host galaxy, with comparable equivalent width, suggesting some host-galaxy extinction.

M. W. Richmond, Rochester Institute of Technology, reports that deep Hubble Space Telescope images of M51 (Hubble Heritage Project) taken in January 2005 show a cluster of young stars near the position of SN 2005cs (cf. IAUC 8553). The object closest to the position of the supernova on Itagaki’s image is at α = 13^{h}29^{m}52^{s}.803, δ = +47^{o}10′36′′.52 (equinox 2000.0). This is a blue star; assuming E(B − V) = 0.10 based on the Na I D absorption reported by Modjaz et al. (above), the candidate has intrinsic colors B-V = −0.2 and V–I = −0.4 and, assuming a distance of 8.4 Mpc to M51, an absolute V magnitude of roughly −6. If this blue star was indeed the progenitor, one might expect SN 2005cs to evolve like SN 1987A, another type-II supernova with a blue progenitor.

G. M. Hurst, Basingstoke, England, reports that an unfiltered CCD image of M51 obtained by P. Birtwhistle (Great Shefford, England, 0.40-m Schmidt-Cassegrain reflector) on June 27.933 UT shows SN 2005cs at red mag 16.6 and at α = 13^{h}29^{m}52^{s}.81, δ = +47^{o}10′35′′.3 (equinox 2000.0).

COMET 161P/HARTLEY-IRAS Total visual magnitude estimates: June 2.10, 11.8 (J. J. Gonzalez,

Leon, Spain, 0.20-m reflector); 3.07, 11.9 (N. Biver, Ablis, France, 0.41-m reflector); 6.09, 11.6 (Gonzalez); 16.08, 11.3 (Gonzalez).

2005 June 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green